# KnowledgeHub: An end-to-end Tool for Assisted Scientific Discovery

### v1

@@ Abstract
This paper describes the KnowledgeHub tool, a scientific literature Information Extraction (IE) and Question Answering (QA) pipeline. This is achieved by supporting the ingestion of PDF documents that are converted to text and structured representations. An ontology can then be constructed where a user defines the types of entities and relationships they want to capture. A browser-based annotation tool enables annotating the contents of the PDF documents according to the ontology. Named Entity Recognition (NER) and Relation Classification (RC) models can be trained on the resulting annotations and can be used to annotate the unannotated portion of the documents. A knowledge graph is constructed from these entity and relation triples which can be queried to obtain insights from the data. Furthermore, we integrate a suite of Large Language Models (LLMs) that can be used for QA and summarisation that is grounded in the included documents via a retrieval component. KnowledgeHub is a unique tool that supports annotation, IE and QA, which gives the user full insight into the knowledge discovery pipeline [^2] A video describing KnowledgeHub is available at this link.
@@ 1 Introduction
Several annotation tools have been created that focus on PDF layout annotation, linguistic annotation or IE through NER and RC. There also exist tools that perform QA over the documents. In the context of the above, tools such as PAWLS Neumann et al. (2021) enable users to annotate PDF document regions with labelled bounding boxes where the tool will then predict layout regions on other files. Linguistic annotation tools such as PDFAnno Shindo et al. (2018), AnnIE Friedrich et al. (2022) and Autodive Du et al. (2023) focus on annotating data for IE tasks such as NER and RC.
Similarly, BatteryDataExtractor Huang and Cole (2022) is a tool that focuses on IE for the battery domain with an additional QA component but does not support annotation.
@@
This paper introduces KnowledgeHub, a novel tool that covers the fundamental aspects of the knowledge discovery process: including linguistic annotation, IE with NER and RC models, and QA that is grounded in the source literature. To achieve this, we implement a pipeline where a user submits a collection of PDF documents for their field of study, which are then converted to text and structured representations. A user-defined ontology can then be supplied which defines the types of entities and relationships to consider. A browser-based annotation tool enables annotating the contents of the PDF documents according to the ontology. NER and RC models can then be trained on the resulting annotations where the trained models can be used to automatically annotate the portion of unannotated documents. A knowledge graph is constructed from these entity and relation triples which can be queried to gain certain insights. Furthermore, we include an RAG based QA system. Out of the systems introduced so far, KnowledgeHub is the only tool that supports annotation, IE, and QA (see Table 1 for an overview of different tools).
@@ 2.1 Document Ingestion
The user uploads an individual PDF file, or a zip file containing multiple PDF files. The PDF content is converted to structured XML using the GROBID tool GRO (2008 2024), which predicts layout sections such as the title, headings, paragraphs, footnotes, references, tables and figures. We extract the text content inside the paragraph objects predicted by GROBID, which is then segmented, tokenised and annotated for Part-of-Speech (POS) information using the Stanza library Qi et al. (2020).
@@ 3 Use-case: Knowledge Discovery for the Battery Domain
Figure 2: Summary of the model answers over retrieved contexts.
Figure 3: The entities/relations contained in the retrieved contexts.

### v2

@@ Abstract
This paper describes the KnowledgeHub tool, a scientific literature Information Extraction (IE) and Question Answering (QA) pipeline. This is achieved by supporting the ingestion of PDF documents that are converted to text and structured representations. An ontology can then be constructed where a user defines the types of entities and relationships they want to capture. A browser-based annotation tool enables annotating the contents of the PDF documents according to the ontology. Named Entity Recognition (NER) and Relation Classification (RC) models can be trained on the resulting annotations and can be used to annotate the unannotated portion of the documents. A knowledge graph is constructed from these entity and relation triples which can be queried to obtain insights from the data. Furthermore, we integrate a suite of Large Language Models (LLMs) that can be used for QA and summarisation that is grounded in the included documents via a retrieval component. KnowledgeHub is a unique tool that supports annotation, IE and QA, which gives the user full insight into the knowledge discovery pipeline.
@@ 1 Introduction
Several annotation tools have been created that focus on PDF layout annotation, linguistic annotation or IE through NER and RC. There also exist tools that perform QA over the documents. In the context of the above, tools such as PAWLS Neumann et al. (2021) enable users to annotate PDF document regions with labelled bounding boxes where the tool will then predict layout regions on other files. Linguistic annotation tools such as PDFAnno Shindo et al. (2018), AnnIE Friedrich et al. (2022) and Autodive Du et al. (2023) focus on annotating data for IE tasks such as NER and RC. Similarly, BatteryDataExtractor Huang and Cole (2022) is a tool that focuses on IE for the battery domain with an additional QA component but does not support annotation.
@@
This paper introduces KnowledgeHub [^2] A video describing KnowledgeHub is available at this link, a novel tool that covers the fundamental aspects of the knowledge discovery process: including linguistic annotation, IE with NER and RC models, and QA that is grounded in the source literature. To achieve this, we implement a pipeline where a user submits a collection of PDF documents for their field of study, which are then converted to text and structured representations. A user-defined ontology can then be supplied which defines the types of entities and relationships to consider. A browser-based annotation tool enables annotating the contents of the PDF documents according to the ontology. NER and RC models can then be trained on the resulting annotations where the trained models can be used to automatically annotate the portion of unannotated documents. A knowledge graph is constructed from these entity and relation triples which can be queried to gain certain insights. Furthermore, we include an RAG based QA system. Out of the systems introduced so far, KnowledgeHub is the only tool that supports annotation, IE, and QA (see Table 1 for an overview of different tools).
@@ 2.1 Document Ingestion
The user uploads an individual PDF file, or a zip file containing multiple PDF files. The PDF content is converted to structured XML using the GROBID tool [^6] https://github.com/kermitt2/grobid, which predicts layout sections such as the title, headings, paragraphs, footnotes, references, tables and figures. We extract the text content inside the paragraph objects predicted by GROBID, which is then segmented, tokenised and annotated for Part-of-Speech (POS) information using the Stanza library Qi et al. (2020).
@@ 3 Use-case: Knowledge Discovery for the Battery Domain
Figure 2: Part of summary of the model answers over retrieved contexts.
Figure 3: Part of the entities/relations contained in the retrieved contexts.